# Multi-Layer Transformers Gradient Can be Approximated in Almost Linear Time

### v1

@@
Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhizhou Sha shazz20@mails.tsinghua.edu.cn. Tsinghua University.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song zsong@adobe.com, magic.linuxkde@gmail.com. Adobe Research.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
The quadratic computational complexity in the self-attention mechanism of popular transformer architectures poses significant challenges for training and inference, particularly in terms of efficiency and memory requirements. Towards addressing these challenges, this paper introduces a novel fast computation method for gradient calculation in multi-layer transformer models. Our approach enables the computation of gradients for the entire multi-layer transformer model in almost linear time $n^{1+o(1)}$, where $n$ is the input sequence length. This breakthrough significantly reduces the computational bottleneck associated with the traditional quadratic time complexity. Our theory holds for any loss function and maintains a bounded approximation error across the entire model. Furthermore, our analysis can hold when the multi-layer transformer model contains many practical sub-modules, such as residual connection, casual mask, and multi-head attention. By improving the efficiency of gradient computation in large language models, we hope that our work will facilitate the more effective training and deployment of long-context language models based on our theoretical results.
@@ 1 Introduction
Large Language Models (LLMs), such as ChatGPT [108], GPT-4 [2], Claude [4], Llama [113, 80], and others, have demonstrated immense potential to enhance various aspects of our daily lives, e.g., conversation AI [61], AI agent [126, 18], search AI [94], AI assistant [93, 139, 53, 32] and many so on. One of the most emergent abilities of LLMs is dealing with long-context information, a format that is crucial for recording material like academic papers, official reports, legal documents, and so on. LLMs have proven adept at tackling long-context tasks, including Retrieval Augmented Generation (RAG) [72, 38], zero-shot summarization [73, 142, 13, 138], and maintaining very long-term conversations [133, 128, 91], and so on. This proficiency has necessitated the development of long-context modeling capabilities within LLMs.
LLMs are mainly based on Transformer architecture, the key module of which is the self-attention mechanism. In attention computation, we will compute the attention score between each pair of tokens, which is the complexity bottleneck during long context training and inference. In detail, we need to spend $O(n^{2}d)$ running time for each self-attention block, which is quadratic in $n$, where $n$ is the length of the context token and $d$ is the hidden feature dimension of the model. For example, LLaMA 3.1 405B [80], one of the cutting-edge LLMs, supports $n=$128k and $d=4096$ taking $30.84$M GPU training hours, which underscores the need for more efficient training processes for such extensive context models. Given the extensive context lengths of LLMs, this quadratic scaling results in several critical challenges: (1) a marked decrease in training efficiency [46, 85, 49]; (2) substantial memory requirements to accommodate the large quadratic gradient matrices [83, 84]; and (3) significant energy usage, which in turn contributes to higher carbon dioxide emissions [110, 102].
One seminar work [5] showed that the self-attention inference can be approximated in almost linear time. However, this result is for the inference time but does not address the main challenge, which is the expensive computation in the training time. In this work, we address this main challenge, by proving that the quadratic computational complexity in the back-propagation of self-attention can be approximated in almost linear time. This suggests we may be able to save the substantial resources required for training Large Language Models.
@@ 1.1 Key background
We first introduce a basic background, starting with defining the softmax function.
@@
Then, we introduce the definition of a single self-attention layer.
Definition 1.2 (Single layer self-attention).
Let $n$ be the number of tokens and $d$ be the hidden dimension size. Let $X\in\mathbb{R}^{n\times d}$ denote the input data matrix. Let $Q:=XW_{Q},K:=XW_{K},V:=XW_{V}\in\mathbb{R}^{n\times d}$. Let $W:=W_{Q}W_{K}^{\top}\in\mathbb{R}^{d\times d}$ be the key-query weight matrix, and let $W_{V}\in\mathbb{R}^{d\times d}$ be the value weight matrix. The self-attention function $\mathsf{Attn}(Q,K,V)$ is:
$\mathsf{Attn}(Q,K,V)=\mathsf{Softmax}(QK^{\top}/d)V.$
We remark that we apply $\mathsf{Softmax}$ to each row of the $n\times n$ matrix.
The above $\mathsf{Attn}$ function takes $Q,K,V$ as parameters, and can be re-parameterized by input matrix $X$ and then re-written in the following sense
$\mathsf{Attn}(X):=f(X)\cdot X\cdot W_{V},$
In contemporary LLMs, the architecture typically incorporates multiple layers of transformers. Consequently, in order to design a fast training algorithm for the entire model, it is imperative to examine the self-attention layers within these multi-layer transformer structures. We provide a formal definition for multi-layer self-attention as follows.
@@
Let $m$ denote the number of transformer layers in the model. Let $X$ be the input sentence. Let $g_{i}$ denote components other than self-attention in the $i$-th transformer layer. Let $\mathsf{Attn}_{i}$ denote the self-attention module in the $i$-th transformer layer (see also Definition 1.2). We define an $m$-layer transformer as
@@
In Definition 1.3, the $g_{i}$ includes the layer norm, MLP, residual connection, dropout, positional encoding, multi-head concatenation, and other operations. All forward and backward computations of these modules can be run in linear time with respect to $n$. Thus, in this work, we mainly focus on the acceleration of self-attention module. Specifically, as shown in Definition 1.2, the $n\times n$ attention matrix $f(X)$ dominates the computational complexity, introducing a quadratic bottleneck. In the exact computation case, if the attention matrix is full rank, no acceleration is possible. However, by compromising negligible accuracy, designing a fast sub-quadratic algorithm becomes feasible. Fortunately, by employing the polynomial kernel approximation method from [1], we can approximate the attention matrix and achieve an almost linear time $n^{1+o(1)}$ algorithm, effectively breaking the quadratic bottleneck.
@@ 1.2 Our contributions
We now state our main contributions as follows:
@@
Let $n$ be the number of tokens, and $d$ be the hidden dimension size. We assume $d=O(\log n)$ and each number in matrices can be written using $O(\log n)$ bits. There exists an algorithm (Algorithm 1) that can compute the gradient of multi-layer self-attention (see also Definition 1.3) in almost linear time $n^{1+o(1)}d=n^{1+o(1)}$, where the approximation error of entire model can be bounded in $1/\operatorname{poly}(n)$.
Our assumption is mild when the context length $n$ is large, as we usually see feature dimension $d$ as a constant, which is also used in [5, 6]. Our results indicate that large language models (LLMs) can be trained in almost linear time $n^{1+o(1)}$ and maintain a robust approximation guarantee, while the traditional way takes $\Omega(n^{2})$ time. This advancement is realized through the application of polynomial kernel approximation [5, 6]. To be more specific, by leveraging the inherent sparsity within the dense attention matrix, we are able to perform efficient low-rank approximation, thereby significantly accelerating the computation of the dense matrices. Our framework is applicable to any loss function, making it universally applicable. Furthermore, our analysis can hold when the multi-layer transformer model contains many practical sub-modules, such as residual connection, casual mask, and multi-head attention (Section 6).
Numerous studies, including Flash Attention [22, 20, 100], quantization techniques [52, 131, 81], and sparsity approaches [44, 54, 88], have empirically focused on accelerating attention mechanisms. However, theoretically, these methods are still constrained by quadratic time complexity. In this study, we introduce an innovative acceleration technique (Algorithm 1) that effectively overcomes this quadratic bottleneck, backed by solid theoretical foundations (Theorem 4.2). Moreover, this new method is designed to be seamlessly integrated with existing approaches to further enhance their performance (see Section 6.4).
@@
  • We introduce a fast computation method that allows the gradient of each self-attention layer to be approximated in almost linear time $n^{1+o(1)}$ with $1/\operatorname{poly}(n)$ error, breaking the quadratic time complexity bottleneck (Theorem 4.1).
  • We extend our single-layer results to module-wise gradient computation so that our Algorithm 1 can provide approximated gradient computation in $m\cdot n^{1+o(1)}$ time for $m$-layer transformer. Furthermore, the approximation of the gradient diverges from the exact gradient by an error of $1/\operatorname{poly}(n)$ when considered across the entire model (Theorem 4.2).
  • Additionally, our analysis can hold when the multi-layer transformer model contains residual connection, casual mask, and multi-head attention. Our results can be applied to any gradient-based algorithm, e.g., training, full fine-tuning, prompt-tuning, and so on (Section 6).
Roadmap. Our paper organized as follows. Section 2 offers the related work. Section 3 provides essential conceptions and key definitions across the whole paper. Section 4 presents our primary findings, where we articulate our novel algorithm that is capable of calculating gradients across the entire model in almost linear time. In Section 5, we explain the techniques we employ, including low-rank approximation, some tricks for accelerating computation of gradients, and an analysis of the approximation error. Section 6 provides various extensions of our algorithm. Lastly, we conclude this paper in Section 7.
@@ 2 Related Work
Attention mechanism.
Attention mechanisms, including self-attention and cross-attention, are pivotal techniques employed in state-of-the-art neural networks. Since it was introduced in [116], it has gained widespread adoption across various domains. In particular, it is integral to decoder-only LLMs [97] and the Vision Transformer (ViT) architecture [21]. The former has been instrumental in the remarkable success of LLMs, while the latter has significantly advanced the field of computer vision, encompassing applications such as image generation [96, 122, 123], detection [70, 141], segmentation [144, 111], and layout generation [34, 19, 118]. Moreover, attention mechanism can be integrated into multi-modal models [127, 137, 77, 121], math reasoning [67], diffusion models [105, 78], differential privacy [9, 104, 117, 76, 99] and many other techniques.
@@
As LLMs grow in size and capability, in-context learning (ICL) [90, 107, 132] has become a preferred method for directing these models to perform a variety of tasks, as opposed to the resource-intensive process of fine-tuning. Nonetheless, research has indicated that longer prompts can impair LLMs performance due to the limitation on maximum sequence length during pre-training [87]. Consequently, extending the maximum sequence length during pre-training and fine-tuning stages is imperative. Enhancing training efficiency is crucial given the prevalent use of the Transformer architecture in LLMs, which incurs a quadratic computational cost relative to sequence length. Addressing this challenge, some studies have explored continued fine-tuning of LLMs with extended context lengths [114], while others have experimented with the interpolation and extrapolation capabilities of positional embedding [17, 95, 98]. However, these approaches have not fundamentally addressed the core issue: the quadratic computational cost associated with sequence length in the attention mechanism [57, 31]. In this study, we delve into accelerating the attention mechanism, thereby addressing the long-context modeling issue at its essence.
@@
Attention mechanism has faced criticism due to its quadratic time complexity with respect to context length, a concern exacerbated by the increasing length in modern large language models (LLMs) such as GPT-4 [2], Claude [4], Llama [113, 80], etc. Nevertheless, this limitation can be circumvented by employing polynomial kernel approximation techniques [1], which enable the derivation of a low-rank representation of the attention matrix. This innovation significantly accelerates both the training and inference processes of a single attention layer, achieving almost linear time complexity [5, 6], while our work supports both training and inference for any multi-layer transformer. Furthermore, this approach can be extended to higher-order attention mechanisms, i.e., tensor attention, maintaining almost linear time complexity during both training and inference [7, 77]. Moreover, there are other theoretical approaches. For instance, [68] introduces the conv-basis method to accelerate attention computation. [44] proposes a near-linear time algorithm under the assumptions of uniform softmax column norms and sparsity.
@@ 3 Preliminary
We first introduce some basic notation used in this paper in Section 3.1. In Section 3.2, we define the loss function. Then, to better understand our proofs, we provide a closed-form analysis of the three gradient components in a single self-attention transformer layer in Section 3.3.
3.1 Notations
For any positive integer $n$, we use $[n]$ to denote set $\{1,2,\cdots,n\}$. For two vectors $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$. Namely, $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. We use $e_{i}$ to denote a vector where only $i$-th coordinate is $1$, and other entries are 00. For each $a,b\in\mathbb{R}^{n}$, we use $a\odot b\in\mathbb{R}^{n}$ to denote the Hardamard product, i.e. the $i$-th entry of $(a\odot b)$ is $a_{i}b_{i}$ for all $i\in[n]$. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $\|A\|_{\infty}$ to denote the $\ell_{\infty}$ norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e. $\|A\|_{\infty}:=\max_{i\in[n],j\in[d]}|A_{i,j}|$. We use $\operatorname{poly}(n)$ to denote polynomial time complexity with respective to $n$.
@@ 3.2 Loss function
For some input matrix $X\in\mathbb{R}^{n\times d}$, we define the one-unit loss function $\ell(X)_{j,k}:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}$, for any $j\in[n],k\in[d]$. Furthermore, we define the overall loss function $L(X)$, such that
@@
Typically, in Definition 3.1, the most widely used loss function in the LLM training procedure is the cross-entropy loss function, which can also be viewed as a summation of one unit loss function.
The output matrix of the multi-layer transformer needs to pass an additional linear layer to map the hidden dimension $d$ to the vocabulary size $d_{\mathrm{voc}}$. Assuming $d_{\mathrm{voc}}$ is a constant, the weight matrix dimensions for this additional MLP layer are $d\times d_{\mathrm{voc}}$. The probability tensor $Y_{\mathrm{pred}}\in\mathbb{R}^{n\times d_{\mathrm{voc}}}$ is the final output. We denote the ground truth as $Y_{\mathrm{gt}}\in\mathbb{R}^{n\times d_{\mathrm{voc}}}$ corresponding to $Y_{\mathrm{pred}}$. According to the cross-entropy loss definition, the formula is expressed as
@@
3.3 Close forms of gradient components

### v2

@@
Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhizhou Sha shazz20@mails.tsinghua.edu.cn. Tsinghua University.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
The computational complexity of the self-attention mechanism in popular transformer architectures poses significant challenges for training and inference, and becomes the bottleneck for long inputs. Is it possible to significantly reduce the quadratic time complexity of computing the gradients in multi-layer transformer models? This paper proves that a novel fast approximation method can calculate the gradients in almost linear time $n^{1+o(1)}$ where $n$ is the input sequence length, while it maintains a polynomially small approximation error $1/\operatorname{poly}(n)$ across the entire model. Our theory holds for general loss functions and when the multi-layer transformer model contains many practical sub-modules, such as residual connection, casual mask, and multi-head attention. By improving the efficiency of gradient computation, we hope that this work will facilitate more effective training and deployment of long-context language models based on our theoretical results.
@@ 1 Introduction
Large Language Models (LLMs), such as ChatGPT [100], GPT-4 [2], Claude 3.5 [4], Llama 3.1 [73], and others, have demonstrated immense potential to enhance various aspects of our daily lives, e.g., conversation AI [49], AI agent [114, 17], search AI [82], AI assistant [81, 124] and many so on. One of the most emergent abilities of LLMs is dealing with long-context information, a format that is crucial for recording material like academic papers, official reports, legal documents, and so on. LLMs have proven adept at tackling long-context tasks, including Retrieval Augmented Generation (RAG) [62, 33], zero-shot summarization [63, 126], and maintaining very long-term conversations [119, 116], and so on. This proficiency has necessitated the development of long-context modeling capabilities within LLMs.
The self-attention mechanism is crucial for the success of LLMs, since LLMs are mainly based on Transformer architecture whose key module is attention. In attention computation, we will compute the attention score between each pair of tokens, which is the complexity bottleneck during long context training and inference. In detail, we need to spend $O(n^{2}d)$ running time for each self-attention block, which is quadratic in $n$, where $n$ is the length of the context input and $d$ is the hidden feature dimension of the model. For example, LLaMA 3.1 405B [73], one of the cutting-edge LLMs, supports $n=$128k and $d=4096$, while taking $30.84$M GPU training hours, which underscores the need for more efficient training processes for such extensive context models. Given the extensive context lengths of LLMs, this quadratic time complexity results in critical challenges: ($i$) a marked decrease in training efficiency [39, 75]; and ($ii$) significant energy usage, which in turn contributes to higher carbon dioxide emissions [102, 91].
One seminal work [5] showed that the self-attention inference can be approximated in almost linear time. However, this result is for the inference time (forward pass), but does not address the main challenge, which is the expensive computation in the training time (backward pass). In this work, we address this main challenge, by proving that the gradient computation in the back-propagation of self-attention can be approximated in almost linear time. This suggests we may be able to save the substantial resources required for training LLMs.
@@ 1.1 Key background
We first introduce some basic background, starting with defining the softmax function and the self-attention module.
@@
Definition 1.2 (Self-attention module).
Let $X\in\mathbb{R}^{n\times d}$ denote the input sequence, where $n$ is the number of input tokens and $d$ is the hidden dimension size. Let $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ be the query, key and value weight matrix. The self-attention function $\mathsf{Attn}(X)$ with weights is:
$\mathsf{Attn}(X)=\mathsf{Softmax}(XW_{Q}W_{K}^{\top}X^{\top}/d)\cdot XW_{V}.$
where $\mathsf{Softmax}$ is applied to each row of its input matrix. The attention can be re-written as:
$\mathsf{Attn}(X)=f(X)\cdot XW_{V},$
where (1) $A:=\exp(XW_{Q}W_{K}^{\top}X^{\top}/d)\in\mathbb{R}^{n\times n}$ and $\exp$ is applied element-wise, (2) $D:=\operatorname{diag}(A{\bf 1}_{n})\in\mathbb{R}^{n\times n}$, and (3) $f(X):=D^{-1}A\in\mathbb{R}^{n\times n}$ is the attention matrix.
In contemporary LLMs, the architecture typically incorporates multiple layers of attention. Consequently, in order to design a fast training algorithm for the entire model, it is imperative to examine self-attention within the multi-layer transformer structure formally defined as follows.
@@
Let $m$ denote the number of transformer layers in the model. Let $X$ be the input sequence. Let $g_{i}$ denote components other than self-attention in the $i$-th transformer layer, and assume its forward and backward computations can be run in time linear in its input sequence length. Let $\mathsf{Attn}_{i}$ denote the self-attention module in the $i$-th transformer layer with weights $W_{Q_{i}},W_{K_{i}},W_{V_{i}}$ (see also Definition 1.2). We define an $m$-layer transformer as
@@
In Definition 1.3, the $g_{i}$ includes the layer norm, MLP, residual connection, dropout, positional encoding, multi-head concatenation, and other operations. All forward and backward computations of these practical modules can be run in linear time with respect to $n$. Thus, in this work, we mainly focus on the acceleration of self-attention module. Specifically, as shown in Definition 1.2, the $n\times n$ attention matrix $f(X)$ dominates the computational complexity, introducing a quadratic bottleneck. In the exact computation case, if the attention matrix is full rank, no acceleration is possible. However, by compromising negligible accuracy, designing a fast sub-quadratic algorithm becomes feasible. Fortunately, by employing the polynomial kernel approximation method from AA [1], we can approximate the attention matrix and achieve an almost linear time $n^{1+o(1)}$ algorithm, effectively breaking the quadratic bottleneck.
@@ 1.2 Our contributions
We now state our main result as follows:
@@
Let $n$ be the number of tokens, and $d$ the hidden dimension size. We assume $d=O(\log n)$ and each number in matrices can be written using $O(\log n)$ bits. Assume the number of layers $m=n^{o(1)}$. There exists an algorithm (Algorithm 1) that can compute the gradient of multi-layer self-attention (see also Definition 1.3) in almost linear time $n^{1+o(1)}$, where the approximation error of the entire model can be bounded by $1/\operatorname{poly}(n)$.
Our assumption is mild when the context length $n$ is large, as the feature dimension $d$ is usually regarded as a constant, which is also used in AA [1]; similarly, the number of layers is usually much smaller than $n$ and regarded as a constant. Our results indicate that large language models (LLMs) can be trained in almost linear time $n^{1+o(1)}$ and maintain a robust approximation guarantee, while the traditional way takes $\Omega(n^{2})$ time. This advancement is realized through the application of polynomial kernel approximation [5, 6]. To be more specific, by leveraging the inherent sparsity within the dense attention matrix, we perform efficient low-rank approximation, thereby significantly accelerating the computation of the dense matrices. Our framework is applicable to general loss functions, making it universally applicable. Furthermore, our analysis holds when the multi-layer transformer model contains many practical sub-modules, such as residual connection, casual mask, and multi-head attention (Section 6).
Numerous studies, including FlashAttention [21, 19, 90], quantization techniques [35, 74], and sparsity approaches [37, 77], have empirically focused on accelerating attention mechanisms. However, theoretically, these methods are still constrained by quadratic time complexity. In this study, we introduce an innovative acceleration technique (Algorithm 1) that effectively overcomes this quadratic bottleneck, backed by solid theoretical foundations (Theorem 4.2). Moreover, this new method is designed to be seamlessly integrated with existing approaches to further enhance their performance (see Section 6).
@@
  • We introduce a fast computation method that allows the gradient of each self-attention layer to be approximated in almost linear time $n^{1+o(1)}$ with $1/\operatorname{poly}(n)$ error, where $n$ is the input sequence length, breaking the quadratic time complexity bottleneck (Theorem 4.1).
  • We extend our single-layer results to module-wise gradient computation so that our Algorithm 1 approximates gradient computation in $m\cdot n^{1+o(1)}$ time for $m$-layer transformer. Importantly, the approximation of the gradient diverges from the exact gradient by an error of $1/\operatorname{poly}(n)$ across the entire model (Theorem 4.2).
  • Additionally, our analysis holds for general loss functions and when the multi-layer transformer model contains residual connection, casual mask, and multi-head attention. Our results can be applied to any gradient-based algorithm, e.g., training, full fine-tuning, prompt-tuning, and so on (Section 6).
@@ 2 Related Work
As LLMs grow in size and capability, in-context learning (ICL) [80, 96, 118, 12] has become a preferred method for directing these models to perform a variety of tasks, as opposed to the resource-intensive process of fine-tuning. Nonetheless, research has indicated that longer prompts can impair LLMs performance due to the limitation on maximum sequence length during pre-training [76]. Consequently, extending the maximum sequence length during pre-training and fine-tuning stages is imperative. Enhancing training efficiency is crucial given the prevalent use of the Transformer architecture in LLMs, which incurs a quadratic computational cost relative to sequence length. Addressing this challenge, some studies have explored continued fine-tuning of LLMs with extended context lengths [104], while others have experimented with the interpolation and extrapolation capabilities of positional embedding [16]. [93] handles long context by compressing the input tokens. However, these approaches have not fundamentally addressed the core issue: the quadratic computational cost associated with sequence length in the attention mechanism [47, 27]. In this study, we delve into accelerating the attention mechanism, thereby addressing the long-context modeling issue at its essence.
@@
Attention mechanism has faced criticism due to its quadratic time complexity with respect to context length, a concern exacerbated by the increasing length in modern large language models (LLMs) such as GPT-4 [2], Claude 3.5 [4], Llama 3.1 [103, 73], etc. Nevertheless, this limitation can be circumvented by employing polynomial kernel approximation techniques [1], which enable the derivation of a low-rank representation of the attention matrix. This innovation significantly accelerates both the training and inference processes of a single attention layer, achieving almost linear time complexity [5, 6], while our work supports both training and inference for any multi-layer transformer. Furthermore, this approach can be extended to higher-order attention mechanisms, i.e., tensor attention, maintaining almost linear time complexity during both training and inference [7, 68]. Moreover, there are other theoretical approaches. For instance, LLS+24d [57] introduces the conv-basis method to accelerate attention computation. HJK+ [37] proposes a near-linear time algorithm under the assumptions of uniform softmax column norms and sparsity.
Roadmap. Our paper is organized as follows. Section 3 provides essential conceptions and key definitions across the whole paper. Section 4 presents our primary findings, where we articulate our novel algorithm that is capable of calculating gradients across the entire model in almost linear time. In Section 5, we explain the techniques we employ, including low-rank approximation, techniques for accelerating the computation of gradients, and an analysis of the approximation error. Section 6 provides various extensions of our algorithm. Lastly, we conclude this paper in Section 7.
@@ 3 Preliminary
Notations.
For any positive integer $n$, we use $[n]$ to denote set $\{1,2,\cdots,n\}$. For two vectors $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$. Namely, $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. We use $e_{i}$ to denote a vector where only $i$-th coordinate is $1$, and other entries are 00. For each $a,b\in\mathbb{R}^{n}$, we use $a\odot b\in\mathbb{R}^{n}$ to denote the Hardamard product, i.e. the $i$-th entry of $(a\odot b)$ is $a_{i}b_{i}$ for all $i\in[n]$. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $\|A\|_{\infty}$ to denote the $\ell_{\infty}$ norm of a matrix $A\in\mathbb{R}^{n\times d}$, i.e., $\|A\|_{\infty}:=\max_{i\in[n],j\in[d]}|A_{i,j}|$. We use $\operatorname{poly}(n)$ to denote some polynomial in $n$.
@@ 3.1 Loss function
For some input matrix $X\in\mathbb{R}^{n\times d}$, we define the one-unit loss function $\ell(X)_{j,k}:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}$, for any $j\in[n],k\in[d]$, and assume differentiability. Furthermore, we define the overall loss function $L(X)$, such that
@@
Typically, the most widely used loss function in the LLM training procedure is the cross-entropy loss function, which can also be viewed as a summation of one unit loss function as in Definition 3.1. The output matrix of the multi-layer transformer needs to pass an additional linear layer to map the hidden dimension $d$ to the vocabulary size $d_{\mathrm{voc}}$. Assuming $d_{\mathrm{voc}}$ is a constant, the weight matrix dimensions for this additional MLP layer are $d\times d_{\mathrm{voc}}$. The probability tensor $Y_{\mathrm{pred}}\in\mathbb{R}^{n\times d_{\mathrm{voc}}}$ is the final output. We denote the ground truth as $Y_{\mathrm{gt}}\in\mathbb{R}^{n\times d_{\mathrm{voc}}}$ corresponding to $Y_{\mathrm{pred}}$. According to the cross-entropy loss definition, the formula is expressed as
@@
3.2 Closed forms of gradient components